OpenStudy (kirbykirby):

\[\int_{\gamma}\frac{z}{(z-1)^{2014}}dz\], where \(\gamma\) is the diamond with vertices \(-2, -2i, 2, 2i \).

OpenStudy (kirbykirby):

I tried using Cauchy's Integral formula: \[f^{(n)}(z_0)=\frac{n!}{2\pi i}\int_{\gamma}\frac{f(z)}{{(z-z_0)}^{n+1}}dz\] So I let \(f(z)=z\) thus giving \(f^{(n)}(z)=0\) and so \(f^{(n)}(-1)=0\) Which implies that the integral must be 0? Although I find it strange because there's the point z=1 inside the closed curve where it's not analytic, and thus this would not satisfy Cauchy's Theorem?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles I'm pretty sure that just because a function isn't analytic on a region with a closed curve \(C\), this doesn't mean that the integral is *not* necessarily zero. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles So just because it doesn't satisfy Cauchy's theorem does not mean that \(\int_C f(z)~dz\not=0\). \(\color{blue}{\text{End of Quote}}\)

OpenStudy (kirbykirby):

Ok good to know!